OpenStudy (anonymous):

Find a number whose product with 6 is the same as its sum with 60. (Hint: you may want to try guess and check or a table.) Show all steps or reasoning. Be sure to check your answer for reasonableness.

OpenStudy (anonymous):

Let the number be = x Its product with 6 will be = 6x Its sum with 60 will be = 60 + x ATQ 6x = 60 + x 6x - x = 60 5x = 60 x = 60/5 x = 12 so the number is 12

OpenStudy (anonymous):

CHECKING: 6x = 6*12 = 72 60 + x = 60 + 12 = 72

OpenStudy (anonymous):

well done harkirat !!!